• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 22:02:16 on November18,2002

    Entry number 90246

    keyword=Hall C crosstalk tests summary

    Assisting Source group with hall C crosstalk studies.

    Parity DAQ, 100uA beam

    Run # time IA (A) IA (C)
    2277 18:58-19:10 5 5 *not a great run b/c no beam for last 6 min b/c when Hall B turned back on, orbit was not good
    2278 19:42-19:52 5 5 *turned hall C current on in middle of run
    2279 19:54-20:00 5 5 *several beam trips and turning hall C beam on when should be off
    2280 20:03-20:12 5 5 *no beam for last 5-6 min of run
    2281 20:31-20:37 5 5 *bcms have been saturating, trying to find good integrate time
    2282 20:37-20:39 5 5 *see previous note
    2283 20:39-20:51 5 5
    2284 20:56-21:02 9 5
    <\PRE>

    MCC called @ 20:12 to see that they had to bypass the cavity, and will be down for at least 20 min.
    Beam back on @ 20:26

    Integrate time (IT) = 13000 for run 2277. Because BCMs were saturating, I lowered
    IT to 12000 for run 2278 & 2279, BCMs still saturating. Lowered IT to 10000 for run
    2280 (Run 2281 had IT=11000). I found that the BCMs were still saturating. Called
    Kent, and set IT=5000, signal then seemed to be about 30000 above pedestal (runs
    2282 &2283), so decided to run with IT=8000 (which is successful with signal around 57000).

    21:45 Tiger laser is having problems. Injector specialist on his way in.
    Called Steve Wood to see if he could come in earlier to start his tests since
    MCC said they could deliver beam to us still. Changed to carbon target

    21:55 MCC called and said North Linac is going into controlled access to fix the Tiger laser. Estimated time = 45 min.